# Temporal Evolution of the Third Interstellar Comet 3I/ATLAS: Spin, Color, Spectra and Dust Activity††thanks: based on observations made with the Southern African Large Telescope (SALT)

### v1

@@ Abstract
Methods. We performed time-series photometry and long-slit spectroscopy between 2 and 29 July 2025 using multiple ground-based telescopes. Photometric data were calibrated against field stars from the ATLAS and APASS catalogs, and Fourier analysis was applied to derive the comet’s rotational period. Spectral data were obtained using the SALT telescope and Nordic Optical Telescope.
@@
Interstellar objects (ISOs) are minor bodies on clearly hyperbolic, Sun-unbound trajectories; their excess speed at infinity confirms an origin beyond the Solar System (Meech et al., 2017). Each ISO offers a physical sample of extrasolar planetesimal material. Studying ISOs is compelling for several reasons: their compositions, activity, and colors probe solid-body formation in other planetary systems; the contrast between the first two ISOs—1I/‘Oumuamua, an apparently inactive, extremely flattened body, and 2I/Borisov, a dust-rich comet—demonstrates the diversity of extrasolar debris reservoirs (Micheli et al., 2018; Jewitt & Luu, 2019); their trajectories reveal how Galactic tides, stellar flybys, and giant-planet scattering shape Oort clouds (Seligman & Moro-Mart´ın, 2023); and their unusual orbits offer rare tests of sublimation and space weathering. Rapid-response ISO campaigns also refine observing strategies ahead of expected ISO discoveries by the Rubin Observatory (Fitzsimmons et al., 2023).
The third confirmed ISO, comet 3I/ATLAS (C/2025 N1), was discovered on 1 July 2025 by the southern ATLAS 0.5 m telescope at Río Hurtado, Chile, and announced as interstellar the next day (Minor Planet Center, 2025). Archival images from ZTF and ATLAS extended the orbit back to mid-June and confirmed an extreme eccentricity, (Seligman et al., 2025). Early follow-up showed that 3I/ATLAS was already active at au, with a dust coma detected by multiple facilities (Bolin et al., 2025; de la Fuente Marcos et al., 2025).
According to JPL Horizons (Giorgini, 2025), 3I/ATLAS will reach perihelion on 29–30 October 2025 at au (just inside Mars’ orbit), then recede into interstellar space. It becomes unobservable from the ground in September due to low solar elongation but re-emerges in the northern sky in late November. Models suggest it will remain visible to ground-based and space telescopes into mid-2026.
In this paper, we present optical photometry and low-resolution spectroscopy of 3I/ATLAS obtained during its first month of observation, while still beyond 3.5 au, establishing a pre-perihelion baseline for tracking its evolution.
@@ 2.1 Spectroscopy
Low-resolution spectroscopic observations of 3I/ATLAS (henceforth 3I) were conducted on 15 and 25 July using the 10-meter Southern African Large Telescope (SALT; see Section A.2.1) and the 2.56-meter Nordic Optical Telescope (NOT; see Section A.2.2), respectively. The SALT spectrum covered the visible wavelength range from to , while for the NOT spectrum we used data from to . To remove the solar continuum and derive the reflectance spectrum of the comet, we observed the solar analogue star SA112-1333 (Landolt, 1992) on the same nights. Spectra of the spectrophotometric standard star Feige 110 (Oke, 1990) were also obtained for flux calibration. Standard data reduction procedures were applied. Individual comet spectra were median-combined and then divided by the solar analogue spectrum. The resulting reflectance spectra are shown in Fig. 1.
@@ 2.2 Photometry
Photometric observations of 3I were conducted between 2 and 29 July 2025 using the Faulkes Telescope North and South (FTN and FTS; see Section A.1.1), the Nordic Optical Telescope (NOT; see Section A.1.2), the Telescopi Joan Oró (TJO; see Section A.1.3), the Lesedi Telescope (see Section A.1.4), and the Skalnaté Pleso Observatory (see Section A.1.5). Time-series photometry was performed exclusively in the SDSS band, while multi-band imaging in and VR filters were used for color measurements. Exposure times were adjusted dynamically based on the seeing conditions and plate scale, maximizing integration time while ensuring stellar point sources remained untrailed. All images were calibrated using standard procedures, including bias subtraction, dark-frame correction, and flat-fielding. Photometric measurements were obtained through aperture photometry relative to selected field stars. Calibration was performed using reference magnitudes from the ATLAS All-Sky Stellar Reference Catalog (Tonry et al., 2018) or from the APASS catalog (Henden & Munari, 2014).
@@ 3.1 Analysis of the color evolution
In the range between – we observed a slight trend of reddening in the comet’s color as it approaches the Sun (see rising and colors in Fig 4), a neutral evolution in the range – (see constant color in Fig 4), and a slight bluing beyond (see decrease in color in Fig 4). However, in general, the measured color remains bluer than reported by Bolin et al. (2025) for 3I at comparable heliocentric distances but our estimates are more consistent with those of Belyakov et al. (2025) and Opitom et al. (2025), who also found relatively neutral to slightly red colors for 3I. When compared to other objects, the color of comet 3I closely resembles that reported for 2I/Borisov (Epifani et al., 2021) at heliocentric distances near 2 au. Moreover, it is similar to distant comets observed at comparable distances around 4 au (e.g., Snodgrass et al. 2008; Ivanova et al. 2015; Voitko et al. 2024), suggesting analogous dust properties among these bodies.
A similar trend is evident in the spectroscopic data, where we derived spectral gradient values from normalized reflectance measurements obtained with NOT and SALT. Based on observations from 4 July, Seligman et al. (2025) reported a gradient of . Our analysis shows an increase to from the SALT data on 15 July, and from NOT observations on 25 July, both calculated over the – range. This progressive reddening over time suggests an evolving surface or coma composition. Beyond , the NOT spectrum becomes noticeably flatter, trending towards a more neutral color. Compared to earlier results by Opitom et al. 2025 and de la Fuente Marcos et al. 2025, we observe a decrease in spectral slope to within –. Future follow-up spectroscopy is planned to track the evolution of the spectral properties, including potential emission lines and ongoing activity indicators.
3.2 Morphology of the coma
@@ 3.3 Magnitude and dust production level
Between 2 and 22 July, photometric observations of 3I revealed a progressive increase in brightness, consistent with its decreasing heliocentric distance from 4.4 au to approximately 3.8 au. The absolute magnitude was determined to be approximately mag, in agreement with the value reported by Jewitt & Luu (2025). Assuming a geometric albedo , our estimate of the comet radius upper limit is about 11 km (see Table 2).
The dust activity level, expressed as , is approximately 300 cm, consistent with estimates by Bolin et al. (2025), but higher than values observed for 2I/Borisov (Guzik et al. 2020; Opitom et al. 2019). This value is comparable to those measured for Jupiter-family comets at large heliocentric distances, as well as for some distant comets exhibiting similar dust activity (Ivanova et al. 2015; Voitko et al. 2024).
Dust mass loss rates are estimated to range from roughly 0.3 to 4.2 kg s-1, slightly higher than those reported by Bolin et al. (2025), but lower than typical rates for Jupiter-family comets at large distances (Gillan et al., 2024), and comparable to distant comets observed at even greater heliocentric distances (Ivanova et al., 2015).
Detailed methodology, equations, and full tables are presented in Appendix B.
@@ 3.4 Lightcurve
Using the time-series photometry described in Section 2.2, we generated a rotational lightcurve of 3I. The comet was observed while crossing the Galactic plane, necessitating mitigation techniques to minimize stellar contamination. We applied Fourier analysis to search for the best-fitting spin period, which was determined to be h (see Figure 3), with a peak-to-peak amplitude of approximately 0.3 mag. The analysis also revealed two prominent alias periods near 8 and 24 hours (see Figure 6).
In some segments of the lightcurve—particularly those corresponding to data from F65 and Z23—the rotational modulation was not clearly detected. These observations often showed a relatively flat lightcurve, lacking distinct maxima or minima and instead clustering around an average magnitude. This behavior is likely due to a combination of observing conditions (e.g., atmospheric seeing, telescope resolution) and the comet’s intrinsic activity, which can enhance the surrounding coma and dilute the nucleus’s photometric variability. In particular, the extensive dataset from Z23 was probably acquired under such conditions, where the rotational signature was partially or entirely suppressed. The inclusion of these flat segments reduced the contrast of the true rotational signal and introduced ambiguity in the Fourier analysis, likely contributing to the emergence of alias solutions.
@@ 4 Conclusions
The morphology of comet 3I during one month of observations reveals clear signs of activity, with continuous dust emission evidenced by the increasing coma size as the heliocentric distance decreased. The absence of a visible tail is likely due to the observing geometry. The coma remained predominantly asymmetric, with no signs of fragmentation or rapid morphological changes. Overall, 3I exhibits characteristics typical of weakly active outer Solar System comets, despite its interstellar origin. Our time-series photometry indicates a spin period of h, with a maximum lightcurve amplitude of 0.3 mag measured at the start of the campaign. This amplitude decreased to 0.2 mag over time, likely as a result of increasing cometary activity, which progressively masked the rotational signal of the nucleus. The presence of enhanced activity is further supported by a reddening trend observed in the comet’s colors throughout the observation period.
We plan to continue photometric and spectroscopic monitoring, and to incorporate polarimetric observations to detect significant changes in color and light-scattering properties of the dust coma in both pre- and post-perihelion data. These observations are essential for characterizing the physical properties of interstellar material, understanding its evolution under solar radiation, and enabling meaningful comparisons with small bodies in the Solar System.
@@ Acknowledgements.
T.S-R. acknowledges funding from Ministerio de Ciencia e Innovación (Spanish Government), PGC2021, PID2021-125883NB-C21. This work was (partially) supported by the Spanish MICIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe” by the “European Union” through grant PID2021-122842OB-C21, and the Institute of Cosmos Sciences University of Barcelona (ICCUB, Unidad de Excelencia ‘María de Maeztu’) through grant CEX2019-000918-M. D.O. was supported by grant No. 2022/45/B/ST9/00267 from the National Science Centre, Poland. O.I. and M.H. thank the Slovak Grants by the Agency for Science, VEGA, grant No. 2/0059/22. This work is based on observations made at the South African Astronomical Observatory (SAAO), which is financially supported by the South African National Research Foundation (NRF). Some of the observations reported in this paper were obtained with the Southern African Large Telescope (SALT). Polish participation in SALT is funded by grant No. MEiN nr 2021/WK/01. The Joan Oró Telescope (TJO) at the Montsec Observatory (OdM) is owned by the Catalan Government and operated by the Institute of Space Studies of Catalonia (IEEC). Partly based on observations made with the Nordic Optical Telescope. The NOT data were obtained under program ID 71-411. Work of Skalnaté Pleso has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. ATLAS is primarily funded to search for near-Earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, and the South African Astronomical Observatory. This paper is based on observations made with the MuSCAT instruments, developed by the Astrobiology Center (ABC) in Japan, the University of Tokyo, and Las Cumbres Observatory (LCOGT). MuSCAT3 was developed with financial support by JSPS KAKENHI (JP18H05439) and JST PRESTO (JPMJPR1775), and is located at the Faulkes Telescope North on Maui, HI (USA), operated by LCOGT. MuSCAT4 was developed with financial support provided by the Heising-Simons Foundation (grant 2022-3611), JST grant number JPMJCR1761, and the ABC in Japan, and is located at the Faulkes Telescope South at Siding Spring Observatory (Australia), operated by LCOGT. Some observations were obtained by the Comet Chasers schools outreach program (https://www.cometchasers.org/), which is funded by the UK Science and Technology Facilities Council (via the DeepSpace2DeepImpact Project), the Open University, and Cardiff University. It accesses the LCOGT telescopes through the Schools Observatory/Faulkes Telescope Project (TSO2025A-00 DFET—The Schools’ Observatory), which is partly funded by the Dill Faulkes Educational Trust, and the LCO Global Sky Partners Programme (LCOEPO2023B-013). Schools making observations included The Coopers Company & Coborn School, Upminster, UK, Ysgol Gyfun Gymraeg Bro Edern, Cardiff, UK, St Marys Catholic Primary School, Bridgend, UK, Institut d’Alcarràs, Catalonia, Spain, Louis Cruis Astronomy Club and Srednja škola Jelkovec (Jelkovec High School), Zagreb, Croatia.References

### v2

@@ Abstract
Methods. We performed time-series photometry and long-slit spectroscopy between 2 and 29 July 2025 using multiple ground-based telescopes. Photometric data were calibrated against field stars from the ATLAS and APASS catalogs, and Fourier analysis was applied to derive the comet’s rotational period. Spectral data were obtained using SALT and Nordic Optical Telescope.
@@
Interstellar objects (ISOs) are minor bodies on hyperbolic, Sun-unbound trajectories, with excess speed at infinity confirming their extrasolar origin (Meech et al., 2017). Each ISO provides a physical sample of extrasolar planetesimals. Studying them is compelling: their compositions, activity, and colors probe solid-body formation in other planetary systems, and the first two ISOs—1I/‘Oumuamua, an apparently inactive, extremely flattened body, and 2I/Borisov, a dust-rich comet—illustrate the diversity of extrasolar debris reservoirs (Micheli et al., 2018; Jewitt & Luu, 2019). Their trajectories reveal how Galactic tides, stellar flybys, and giant-planet scattering shape Oort clouds (Seligman & Moro-Mart´ın, 2023), while unusual orbits offer rare tests of sublimation and space weathering. Rapid-response ISO campaigns also refine strategies for anticipated Rubin Observatory discoveries (Fitzsimmons et al., 2023).
The third confirmed ISO, comet 3I/ATLAS (C/2025 N1), was discovered on 1 July 2025 by the ATLAS 0.5 m telescope at Río Hurtado, Chile, and announced as interstellar the next day (Minor Planet Center, 2025). Archival ZTF and ATLAS images extended the orbit to mid-June, confirming extreme eccentricity, (Seligman et al., 2025). Early follow-up showed 3I/ATLAS was already active at au, with a dust coma detected by multiple facilities (Bolin et al., 2025; de la Fuente Marcos et al., 2025); TESS (Feinstein et al., 2025; Martinez-Palomera et al., 2025) and Rubin/LSST (Chandler et al., 2025) precovery data further extend the observational arc.
JPL Horizons (Giorgini, 2025) predicts perihelion on 29–30 October 2025 at au, just inside Mars’ orbit, before it recedes into interstellar space. Ground-based observations will be impossible from September due to low solar elongation, with the object reappearing in the northern sky in late November and remaining visible until mid-2026.
Here, we present optical photometry and low-resolution spectroscopy of 3I/ATLAS obtained during its first month of observations, still beyond 3.5 au, establishing a pre-perihelion baseline for tracking its evolution.
@@ 2.1 Spectroscopy
Low-resolution spectroscopic observations of 3I/ATLAS (hereafter 3I) were obtained on 15 and 25 July using the 10-meter Southern African Large Telescope (SALT; see Section A.2.1) and the 2.56-meter Nordic Optical Telescope (NOT; see Section A.2.2), respectively. The SALT spectrum spanned –, while the NOT spectrum covered –. To remove the solar continuum and derive the comet’s reflectance spectrum, we observed the solar analogue SA112-1333 (Landolt, 1992) on the same nights. Spectra of the spectrophotometric standard Feige 110 (Oke, 1990) were also obtained for flux calibration. Standard data reduction procedures were applied, after which individual comet spectra were median-combined and divided by the solar analogue. The resulting reflectance spectra are shown in Fig. 1.
@@ 2.2 Photometry
Photometric observations of 3I were carried out between 2 and 29 July 2025 using the Faulkes Telescopes North and South (FTN and FTS; see Section A.1.1), the Nordic Optical Telescope (NOT; see Section A.1.2), the Telescopi Joan Oró (TJO; see Section A.1.3), the Lesedi Telescope (see Section A.1.4), and the Skalnaté Pleso Observatory (see Section A.1.5). Time-series photometry was performed in the SDSS band, while multi-band imaging in and VR filters provided color measurements. Exposure times were adjusted dynamically according to seeing and plate scale, maximizing integration while keeping stellar point sources untrailed. All images were calibrated with standard procedures, including bias subtraction, dark-frame correction, and flat-fielding. Photometry was obtained via aperture photometry relative to selected field stars, with calibration based on reference magnitudes from the ATLAS All-Sky Stellar Reference Catalog (Tonry et al., 2018) or the APASS catalog (Henden & Munari, 2014).
@@ 3.1 Analysis of the color evolution
In the – range we observed slight reddening as the comet approached the Sun (rising and in Fig. 2), a neutral trend at – (constant ), and slight bluing beyond (decreasing ). Overall, the color is bluer than reported by Bolin et al. (2025) for 3I at similar distances, but closer to Belyakov et al. (2025) and Opitom et al. (2025), who also found neutral to slightly red values. Compared to other bodies, 3I’s color resembles that of 2I/Borisov near 2 au (Epifani et al., 2021) and distant comets near 4 au (e.g., Snodgrass et al. 2008; Ivanova et al. 2015; Voitko et al. 2024), suggesting analogous dust properties.
Spectroscopy shows a similar trend: spectral gradients from NOT and SALT reflectance data indicate progressive reddening. Seligman et al. (2025) reported on 4 July, while we measure (SALT, 15 July) and (NOT, 25 July) over –. Beyond , the NOT spectrum flattens to at –, consistent with Opitom et al. 2025 and de la Fuente Marcos et al. 2025. In contrast, NIR spectra (Kareta et al., 2025; Yang et al., 2025) show strong flattening beyond m, with slopes , implying a transition to neutral or slightly blue slopes. Thus, 3I shows wavelength-dependent evolution: redder in the optical, flatter or bluer in the NIR, consistent with the color trend. Future spectroscopy will track slope changes, emission lines, and activity.
@@ 3.2 Magnitude and dust production level
Between 2 and 29 July, photometric observations of 3I show a progressive brightening, consistent with its decreasing heliocentric distance from 4.5 au to au. The absolute magnitude was mag, in agreement with the reported by Jewitt & Luu (2025). Assuming a geometric albedo , the estimated upper limit on the comet radius is km (see Table 2). The dust activity level, expressed as , is cm, consistent with Bolin et al. (2025), but higher than values observed for 2I/Borisov (Guzik et al. 2020; Opitom et al. 2019). This is comparable to measurements of Jupiter-family comets at large heliocentric distances and some distant comets exhibiting similar activity (Ivanova et al. 2015; Voitko et al. 2024). Estimated dust mass-loss rates range from to kg s-1, slightly above those reported by Bolin et al. (2025), yet lower than typical rates for Jupiter-family comets at large distances (Gillan et al., 2024), and comparable to distant comets observed even farther from the Sun (Ivanova et al., 2015). Detailed methodology, relevant equations, and complete tables are provided in Appendix C.
@@ 3.3 Lightcurve
Using the time-series photometry described in Section 2.2, we generated a rotational lightcurve of 3I. Observations occurred while the comet crossed the Galactic plane, requiring mitigation techniques to minimize stellar contamination. Fourier analysis yielded a best-fitting spin period of h (see Figure 3), with a peak-to-peak amplitude of mag and prominent aliases near 8 and 24 h (see Figure 5). This broadly agrees with the h period reported by de la Fuente Marcos et al. (2025), though their estimate—based on 2–5 July photometry—may be affected by the limited temporal baseline.
The lightcurve amplitude can be related to the projected axis ratio of a triaxial ellipsoid via (e.g., Binzel & Sauter, 1989), giving –0.83 for –0.3 mag, which represents minimum elongations for equatorial viewing and small phase angles (–).
Using the Hubble Space Telescope (HST) upper limit on the nucleus radius, km (Jewitt et al., 2025), the maximum principal axes are
@@ 4 Conclusions
The morphology of comet 3I over one month reveals clear activity, with continuous dust emission evidenced by the increasing coma size as the heliocentric distance decreased. The absence of a visible tail likely reflects the observing geometry. The coma remained predominantly asymmetric, with no signs of fragmentation or rapid changes. Overall, 3I shows characteristics typical of weakly active outer Solar System comets, despite its interstellar origin. Time-series photometry indicates a spin period of h, with a maximum lightcurve amplitude of 0.3 mag at the start, decreasing to 0.2 mag—likely as increasing activity progressively masks the nucleus’s rotational signal. The lightcurve amplitude relates to the projected axis ratio via (e.g., Binzel & Sauter, 1989), giving –0.83 for –0.3 mag, representing minimum elongations for equatorial viewing and small phase angles (–). Enhanced activity is further supported by a reddening trend in the comet’s colors.
Measured values (300 cm) are comparable to distant Jupiter-family comets (Gillan et al., 2024), but dust production rates are lower than expected for an object of this brightness at au, suggesting a coma dominated by large particles. Independent HST observations report dust mass-loss rates of 6–60 kg s-1 (Jewitt et al., 2025), also indicating large grains. The absence of a distinct tail and the homogeneous coma morphology support this, as massive particles are less affected by solar radiation pressure and remain near the nucleus, implying a deficit of submicron dust that would otherwise enhance light scattering. Further polarimetric and infrared observations are needed to constrain the particle size distribution. We plan to continue photometric and spectroscopic monitoring, including polarimetry to detect changes in color and light-scattering properties of the coma pre- and post-perihelion. These observations are essential to characterize interstellar material, its evolution under solar radiation, and to enable comparisons with Solar System small bodies.
@@ Acknowledgements.
T.S-R. acknowledges funding from Ministerio de Ciencia e Innovación (Spanish Government), PGC2021, PID2021-125883NB-C21. This work was (partially) supported by the Spanish MICIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe” by the “European Union” through grant PID2021-122842OB-C21, and the Institute of Cosmos Sciences University of Barcelona (ICCUB, Unidad de Excelencia ‘María de Maeztu’) through grant CEX2019-000918-M. D.O. was supported by grant No. 2022/45/B/ST9/00267 from the National Science Centre, Poland. O.I. and M.H. thank the Slovak Grants by the Agency for Science, VEGA, grant No. 2/0059/22. This work is based on observations made at the South African Astronomical Observatory (SAAO), which is financially supported by the South African National Research Foundation (NRF). Some of the observations reported in this paper were obtained with the Southern African Large Telescope (SALT). Polish participation in SALT is funded by grant No. MEiN nr 2021/WK/01. The Joan Oró Telescope (TJO) at the Montsec Observatory (OdM) is owned by the Catalan Government and operated by the Institute of Space Studies of Catalonia (IEEC). Partly based on observations made with the Nordic Optical Telescope. The NOT data were obtained under program ID 71-411 (PI A. Penttilä). Work of Skalnaté Pleso has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. ATLAS is primarily funded to search for near-Earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, and the South African Astronomical Observatory. This paper is based on observations made with the MuSCAT instruments, developed by the Astrobiology Center (ABC) in Japan, the University of Tokyo, and Las Cumbres Observatory (LCOGT). MuSCAT3 was developed with financial support by JSPS KAKENHI (JP18H05439) and JST PRESTO (JPMJPR1775), and is located at the Faulkes Telescope North on Maui, HI (USA), operated by LCOGT. MuSCAT4 was developed with financial support provided by the Heising-Simons Foundation (grant 2022-3611), JST grant number JPMJCR1761, and the ABC in Japan, and is located at the Faulkes Telescope South at Siding Spring Observatory (Australia), operated by LCOGT. Some observations were obtained by the Comet Chasers schools outreach program (https://www.cometchasers.org/), which is funded by the UK Science and Technology Facilities Council (via the DeepSpace2DeepImpact Project), the Open University, and Cardiff University. It accesses the LCOGT telescopes through the Schools Observatory/Faulkes Telescope Project (TSO2025A-00 DFET—The Schools’ Observatory), which is partly funded by the Dill Faulkes Educational Trust, and the LCO Global Sky Partners Programme (LCOEPO2023B-013). Schools making observations included The Coopers Company & Coborn School, Upminster, UK, Ysgol Gyfun Gymraeg Bro Edern, Cardiff, UK, St Marys Catholic Primary School, Bridgend, UK, Institut d’Alcarràs, Catalonia, Spain, Louis Cruis Astronomy Club and Srednja škola Jelkovec (Jelkovec High School), Zagreb, Croatia.References
@@
- Binzel & Sauter (1989) Binzel, R. P., & Sauter, L. M. 1989, in Asteroids II, Univ. of Arizona Press, 416–441
@@
- Chandler et al. (2025) Chandler, C. O., Bernardinelli, P. H., Jurić, M., et al. 2025, NSF-DOE Vera C. Rubin Observatory Observations of Interstellar Comet 3I/ATLAS (C/2025 N1), arXiv:2507.13409, https://arxiv.org/abs/2507.13409
@@
- Feinstein et al. (2025) Feinstein, A. D., Noonan, J. W., & Seligman, D. Z. 2025, Precovery Observations of 3I/ATLAS from TESS Suggest Possible Distant Activity, arXiv:2507.21967, https://arxiv.org/abs/2507.21967
@@
- Ivanova et al. (2019) Ivanova, O., Luk’yanyk, I., Kolokolova, L., et al. 2019, A&A, 626, A26
- Ivanova et al. (2021) Ivanova, O., Rosenbush, V., Luk’yanyk, I., et al. 2021, A&A, 651, A29
- Ivanova et al. (2023) Ivanova, O., Rosenbush, V., Luk’yanyk, I., et al. 2023, A&A, 672, A76
@@
- Jewitt et al. (2025) Jewitt, D., Hui, M.-T., Mutchler, M., Kim, Y., & Agarwal, J. 2025, Hubble Space Telescope Observations of the Interstellar Interloper 3I/ATLAS, arXiv:2508.02934, https://arxiv.org/abs/2508.02934
- Kareta et al. (2025) Kareta, T., Champagne, C., McClure, L., et al. 2025, arXiv e-print arXiv:2507.12234
@@
- Manzini et al. (2020) Manzini, F., Oldani, V., Ochner, P., & Bedin, L. R. 2020, MNRAS Lett., 495, L92
- Martinez-Palomera et al. (2025) Martinez-Palomera, J., Tuson, A., Hedges, C., et al. 2025, Pre-discovery TESS Observations of Interstellar Object 3I/ATLAS, arXiv:2508.02499, https://arxiv.org/abs/2508.02499
@@
- Meech & Svoren (2004) Meech, K. J., & Svoren, J. 2004, in Comets II, ed. M. C. Festou, H. U. Keller, & H. A. Weaver (Tucson: Univ. of Arizona Press), 317
@@
- Yang et al. (2025) Yang, B., Meech, K. J., Connelley, M., & Keane, J. V. 2025, arXiv e-print arXiv:2507.14916